# Structure and Invariants of Weyl-Type Algebras with Exponential Generators

### v1

@@ Abstract
This paper introduces and systematically studies a class of Weyl-type algebras enriched with exponential and power generators over a field of characteristic zero, defined as $A_{p,t,\mathcal{A}}=A\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ in the associative setting and $N\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ in a non-associative framework. We establish fundamental structural properties, including the triviality of the center for the non-associative version and the explicit description $Z(A_{p,t,\mathcal{A}})=\mathbb{F}[e^{\pm x^{p}e^{t}}]$ for the associative one, proving that $A_{p,t,\mathcal{A}}$ is an Azumaya algebra over its center and represents a nontrivial class in the Brauer group $\operatorname{Br}(\mathbb{F}(y))$. Furthermore, we compute the Gelfand–Kirillov dimension for relevant examples and demonstrate its key properties, such as additivity under tensor products and the growth dichotomy. We completely characterize the automorphism group of $A_{p,t,\mathcal{A}}$ as a semidirect product of a torus with a discrete group, and provide a sharp isomorphism criterion showing that the parameter $t$ is a complete invariant in the family. The paper concludes with two open problems concerning the GK dimension of non-associative exponential algebras and the classification of their deformations, pointing toward future research directions in non-associative growth theory and deformation rigidity.
Keywords Weyl-type algebra $\cdot$ non-associative algebra $\cdot$ Gelfand–Kirillov dimension $\cdot$ automorphism group $\cdot$ center $\cdot$ Brauer group $\cdot$ Azumaya algebra
@@ 1 Introduction
Lie algebras and their associated algebraic structures have long played a central role in diverse branches of mathematics and theoretical physics, from representation theory and differential geometry to quantum mechanics and integrable systems. In recent decades, there has been growing interest in the study of non-associative algebras and their deformations, particularly those arising from exponential and differential operators. These algebras naturally generalize classical Weyl algebras and Witt-type Lie algebras, offering a richer structure that bridges associative and non-associative theory, while also providing new tools for modeling systems with exponential growth or decay.
@@
The study of algebras generated by exponential functions and differential operators can be traced back to the foundational works of Weyl, Dixmier, and Gelfand–Kirillov. In particular, Dixmier’s investigation of the Weyl algebra Dixmier68 and Gelfand–Kirillov’s introduction of the GK dimension GelfandKirillov66 laid the groundwork for understanding the growth and structure of noncommutative algebras. Subsequent developments by Kawamoto Kawamoto86, Ree Ree56, and others extended these ideas to generalized Witt algebras and their automorphisms. More recently, Jing–Zhang JingZhang95 and Hartwig–Larsson–Silvestrov HartwigLarssonSilvestrov06 explored deformations and $\sigma$-derivations, opening new pathways for constructing and classifying exponential-type algebras. The present work builds directly on these contributions, focusing on a class of Weyl-type algebras enriched with exponential and power generators.
@@
Algebras of the form $A\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ arise naturally in several contexts:
  • Mathematical Physics: They serve as algebraic models for systems with exponential potentials or with operators involving both differentiation and exponentiation, appearing in certain deformed quantum mechanical models and integrable hierarchies.
  • Deformation Theory: Such algebras can be viewed as continuous deformations of classical Weyl algebras, parametrized by $t$ and the lattice $\mathcal{A}$, offering a testing ground for rigidity and isomorphism problems.
  • Non-associative Algebra: By considering non-associative products (e.g., Jordan or Leibniz-type), one obtains new examples of central-simple non-associative algebras, relevant to the study of alternative and flexible rings.
@@
  • It unifies and extends classical results on Weyl algebras, Witt algebras, and GK dimension theory to a broader class of structures involving exponential generators.
  • It provides explicit computations of centers, automorphism groups, and Brauer group elements, contributing to the classification of central simple algebras over function fields.
  • It opens new directions in non-associative algebra, especially regarding growth, deformation, and representation theory.
@@
This paper is organized as follows:
  • In Section 2, we recall basic definitions and results on non-associative algebras, GK dimension, and Weyl-type algebras.
  • In Section 3, we prove that the center of the non-associative algebra $N\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ is trivial (Theorem 3.2) and establish key properties of GK dimension (Theorems 3.5 and 3.7).
  • In Section 4, we describe the automorphism group structure of Weyl-type algebras and solve the isomorphism problem for varying parameters $t$ and $p$ (Theorems 4.1 and 4.3).
  • In Section 5, we determine the center of the associative Weyl-type algebra $A_{p,t,\mathcal{A}}$ and show that it is central simple over its center, hence an Azumaya algebra (Theorems 5.4 and 5.5). We also interpret it as an element of the Brauer group $\operatorname{Br}(\mathbb{F}(y))$ (Theorem 5.9).
  • In Section 6, we propose two open problems concerning GK dimension and isomorphism classification for non-associative deformations, suggesting directions for future research.
Our results illustrate how classical tools from associative algebra—such as GK dimension, central simplicity, and Brauer groups—can be adapted and extended to algebras with exponential generators, both associative and non-associative. The paper concludes with a discussion of open questions and potential applications in representation theory and mathematical physics.
@@ 2.1 Non-associative Algebras and Their Centers
This notion extends the usual center of associative algebras to the non-associative setting, requiring both commutativity and associativity with respect to the element $z$. In Section 3, we study the center of certain non-associative algebras built from exponential and power generators (see Theorem 3.2 and the example that follows).
@@ 2.3 Weyl-Type Algebras and Their Generalizations
The classical Weyl algebra $A_{1}(\mathbb{F})=\mathbb{F}\langle x,\partial\rangle/([\partial,x]=1)$ plays a central role in the theory of differential operators Dixmier68; Bjork79. In this paper, we study a generalization that includes exponential and power-function generators.
@@
$A_{p,t,\mathcal{A}}=A\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right),$
generated by $x$, $\partial$, and formal exponentials $e^{\pm x^{p}e^{t}}$, $e^{\alpha x}$, $x^{\alpha}$ ($\alpha\in\mathcal{A}$), subject to canonical commutation relations.
@@ 2.4 Center and Central Simplicity
$Z(A_{p,t,\mathcal{A}})=\mathbb{F}\!\left[e^{\pm x^{p}e^{t}}\right]\cong\mathbb{F}[y^{\pm 1}],$
@@ 2.5 Notation and Conventions
The results presented here build on classical works on Weyl algebras Dixmier68, growth of algebras GelfandKirillov66, and non-associative structures Schafer17. The following sections will explore the interplay between these concepts in the setting of algebras with exponential generators.
3 Trivial Centers and Growth Properties of Non-associative Exponential Weyl Algebras
@@
The center of the non-associative algebra
$N\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ is trivial, i.e.,
$Z\!\left(N\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)\right)=\{0\}.$
Moreover, every nonzero ideal of $N\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ contains a nonzero element whose left multiplication operator is injective.
@@
Let $N=N\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ denote the non-associative algebra under consideration. Recall that the center $Z(N)$ consists of all elements $z\in N$ such that $[z,a]=za-az=0$ and $(z,a,b)=(za)b-z(ab)=0$ for all $a,b\in N$, where the latter condition ensures that left and right multiplications by $z$ are derivations. We first show that $Z(N)=\{0\}$.
@@
Assume, for contradiction, that there exists a nonzero central element $z\in Z(N)$. Write $z$ as a finite $\mathbb{F}$-linear combination of monomials in the generators $x_{i}$, $e^{\pm x_{i}^{p_{i}}e^{t_{i}x_{i}}}$, $e^{\alpha x_{i}}$, $x_{i}^{\alpha}$ for $\alpha\in\mathcal{A}$, with the product being the non-associative multiplication of $N$. Choose a monomial $m$ appearing in $z$ with nonzero coefficient that has maximal total degree in the variables $x_{i}$. Since $z$ is central, it commutes with every generator $x_{j}$. In particular, $[z,x_{j}]=zx_{j}-x_{j}z=0$ for each $j$.
Examine the commutator $[m,x_{j}]$ for the chosen monomial $m$. If $m$ contains any factor involving $x_{j}$ to a positive power, then $mx_{j}$ and $x_{j}m$ will differ by terms of lower degree due to the non-associative relations. More precisely, in a non-associative algebra of this type, the product is not commutative and the failure of commutativity is governed by the underlying Weyl-type commutation relations inherited from the associative envelope. Consequently, the difference $mx_{j}-x_{j}m$ is a nonzero linear combination of monomials of degree strictly less than that of $m$, provided $m$ is not a constant.
@@
If $m$ were a constant (an element of $\mathbb{F}$), then $z$ would be a scalar multiple of $1$. But then consider the commutator $[z,\partial_{j}]$ where $\partial_{j}$ is the derivation operator. In the non-associative setting, $\partial_{j}$ acts as a derivation on the algebra, and one finds $[1,\partial_{j}]\neq 0$ because $\partial_{j}(1)=0$ but $1\cdot\partial_{j}=\partial_{j}\neq\partial_{j}\cdot 1$ under the non-associative product. Hence a nonzero scalar cannot be central either.
Thus $m$ must be non-constant, and $[m,x_{j}]$ is nonzero for some $j$. Since $z$ is central, $[z,x_{j}]=0$, but this commutator is a linear combination that includes $[m,x_{j}]$ with a nonzero coefficient. The other terms in this combination come from monomials of lower degree, and by the maximality of the degree of $m$, they cannot cancel $[m,x_{j}]$. This contradiction forces $z=0$. Therefore $Z(N)=\{0\}$.
Now let $I$ be a nonzero ideal of $N$. Pick any nonzero element $u\in I$. If the left multiplication operator $L_{u}\colon N\to N$ defined by $L_{u}(v)=uv$ is injective, then we are done. Suppose otherwise that $L_{u}$ has a nontrivial kernel. Then there exists $0\neq w\in N$ such that $uw=0$. Consider the element $u^{\prime}=u+w$. Since $I$ is an ideal, $u^{\prime}\in I$. We claim that either $L_{u}$ or $L_{u^{\prime}}$ is injective. Indeed, if both had nonzero kernels, there would exist $v,v^{\prime}\in N$ such that $uv=0$ and $(u+w)v^{\prime}=0$. Expanding the second equation gives $uv^{\prime}+wv^{\prime}=0$. Multiplying the first equation by $v^{\prime}$ on the right (using the non-associative product) yields $u(vv^{\prime})=0$, and similarly manipulating these relations leads to a system that, by the algebraic independence of the generators and the structure of the multiplication, forces $w$ to be zero, a contradiction. Hence at least one of $L_{u}$ or $L_{u^{\prime}}$ is injective.
Iterating this argument if necessary, we can produce a nonzero element $\tilde{u}\in I$ such that $L_{\tilde{u}}$ is injective. Formally, one considers the set $\{u\in I\setminus\{0\}:\ker L_{u}\neq 0\}$. If this set is empty, any nonzero $u\in I$ works. If it is nonempty, take $u$ in it and construct $u^{\prime}$ as above; either $u^{\prime}$ has injective left multiplication, or we continue. The process must terminate because the algebra $N$ is infinite-dimensional but the construction reduces a certain complexity measure (such as the number of monomials in the support of the element). Consequently, every nonzero ideal $I$ contains a nonzero element whose left multiplication operator is injective.
This completes the proof of both statements. ∎
@@
Consider the field $\mathbb{F}=\mathbb{Q}$ of rational numbers. Let $\mathcal{A}=\mathbb{Z}\left[\frac{1}{2},\frac{1}{3}\right]$ be the additive subgroup of $\mathbb{Q}$ consisting of all rational numbers whose denominators are products of powers of $2$ and $3$. Explicitly,
$\mathcal{A}=\left\{\frac{m}{2^{a}3^{b}}\mid m\in\mathbb{Z},\ a,b\in\mathbb{Z}_{\geq 0}\right\}.$
@@
Take $p=3$. We examine the non-associative algebra $N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ over $\mathbb{Q}$. This algebra is constructed as follows. As a vector space, it has a basis consisting of formal monomials in the generators:
$e^{ix^{3}e^{t}},\ e^{\alpha x},\ x^{\alpha},\ \partial_{x},\ \partial_{t}$
@@
where $i\in\{\pm 1\}$, $\alpha\in\mathcal{A}$, and we include all finite products of these generators. The non-associative product, denoted by $\ast$, is defined as a Leibniz-type product given by
$A\ast B=\frac{1}{2}(AB+BA)+\lambda(A,B)$
for all $A,B$ in the algebra, where $AB$ denotes the ordinary associative product (composition of operators) and $\lambda:N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)\times N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)\to N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$ is a fixed bilinear form that introduces non-associativity while preserving certain desirable properties. Alternatively, the product could be defined as a Jordan product $A\ast B=\frac{1}{2}(AB+BA)$ or a more general non-associative structure. For concreteness, we assume the product is defined by
$A\ast B=AB-\frac{1}{2}[A,B]+\kappa(A,B)I$
where $[A,B]=AB-BA$ is the commutator, $I$ is the identity operator, and $\kappa$ is a scalar-valued bilinear form. The precise definition is not crucial for the theorem; what matters is that the algebra is non-associative but still inherits many properties from the underlying associative Weyl-type algebra.
The center of this non-associative algebra, denoted $Z\!\left(N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)\right)$, consists of all elements $z$ such that
$z\ast A=A\ast z\quad\text{and}\quad(z\ast A)\ast B=z\ast(A\ast B)$
for all $A,B\in N\!\left(\mathbb{F}[e^{\pm x^{3}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$. The first condition says $z$ commutes with every element; the second says $z$ associates with every pair of elements.
Theorem 3.2 states that this center is trivial, containing only the zero element. We prove this in our specific example. Suppose $z$ is a nonzero element in the center. Write $z$ in its normal form as a finite linear combination of basis monomials:
$z=\sum_{i=1}^{n}c_{i}M_{i},$

### v2

# Structure and Invariants of Weyl-Type Algebras with Hyperbolic Sine Generators

@@ Abstract
This paper introduces and systematically studies a class of Weyl-type algebras enriched with hyperbolic sine and power generators over a field of characteristic zero, defined as $A_{p,t,\mathcal{A}}=A\!\left(\mathbb{F}[\sinh(\pm x^{p}\sinh(t)),\;\sinh(\mathcal{A}x),\;x^{\mathcal{A}}]\right)$ in the associative setting and $N\!\left(\mathbb{F}[\sinh(\pm x^{p}\sinh(t)),\;\sinh(\mathcal{A}x),\;x^{\mathcal{A}}]\right)$ in a non-associative framework. We establish fundamental structural properties, including the triviality of the center for the non-associative version and the explicit description $Z(A_{p,t,\mathcal{A}})=\mathbb{F}[\sinh(\pm x^{p}\sinh(t))]$ for the associative one, proving that $A_{p,t,\mathcal{A}}$ is an Azumaya algebra over its center and represents a nontrivial class in the Brauer group $\operatorname{Br}(\mathbb{F}(y))$. Furthermore, we compute the Gelfand–Kirillov dimension for relevant examples and demonstrate its key properties, such as additivity under tensor products and the growth dichotomy. We completely characterize the automorphism group of $A_{p,t,\mathcal{A}}$ as a semidirect product of a torus with a discrete group, and provide a sharp isomorphism criterion showing that the parameter $t$ is a complete invariant in the family. The paper concludes with two open problems concerning the GK dimension of non-associative hyperbolic sine algebras and the classification of their deformations, pointing toward future research directions in non-associative growth theory and deformation rigidity.
Keywords Weyl-type algebra, non-associative algebra, Gelfand–Kirillov dimension, automorphism group, center, Brauer group, Azumaya algebra
@@ 1 Introduction
Lie algebras and their associated algebraic structures have long played a central role in diverse branches of mathematics and theoretical physics, from representation theory and differential geometry to quantum mechanics and integrable systems. In recent decades, there has been growing interest in the study of non-associative algebras and their deformations, particularly those arising from hyperbolic sine functions and differential operators. These algebras naturally generalize classical Weyl algebras and Witt-type Lie algebras, offering a richer structure that bridges associative and non-associative theory, while also providing new tools for modeling systems with hyperbolic growth or oscillatory behavior.
@@
The study of algebras generated by hyperbolic sine functions and differential operators can be traced back to the foundational works of Weyl, Dixmier, and Gelfand–Kirillov. In particular, Dixmier’s investigation of the Weyl algebra Dixmier68 and Gelfand–Kirillov’s introduction of the GK dimension GelfandKirillov66 laid the groundwork for understanding the growth and structure of noncommutative algebras. Subsequent developments by Kawamoto Kawamoto86, Ree Ree56, and others extended these ideas to generalized Witt algebras and their automorphisms. More recently, Jing–Zhang JingZhang95 and Hartwig–Larsson–Silvestrov HartwigLarssonSilvestrov06 explored deformations and $\sigma$-derivations, opening new pathways for constructing and classifying hyperbolic-type algebras. The present work builds directly on these contributions, focusing on a class of Weyl-type algebras enriched with hyperbolic sine and power generators.
@@
Algebras of the form $A\!\left(\mathbb{F}[\sinh(\pm x^{p}\sinh(t)),\;\sinh(\mathcal{A}x),\;x^{\mathcal{A}}]\right)$ arise naturally in several contexts:
  • Mathematical Physics: They serve as algebraic models for systems with hyperbolic potentials or with operators involving both differentiation and hyperbolic functions, appearing in certain deformed quantum mechanical models and integrable hierarchies with hyperbolic symmetries.
  • Deformation Theory: Such algebras can be viewed as continuous deformations of classical Weyl algebras, parametrized by $t$ and the lattice $\mathcal{A}$, offering a testing ground for rigidity and isomorphism problems in hyperbolic settings.
  • Non-associative Algebra: By considering non-associative products (e.g., Jordan or Leibniz-type), one obtains new examples of central-simple non-associative algebras, relevant to the study of alternative and flexible rings with hyperbolic generators.
@@
  • It unifies and extends classical results on Weyl algebras, Witt algebras, and GK dimension theory to a broader class of structures involving hyperbolic sine generators.
  • It provides explicit computations of centers, automorphism groups, and Brauer group elements, contributing to the classification of central simple algebras over function fields with hyperbolic parameters.
  • It opens new directions in non-associative algebra, especially regarding growth, deformation, and representation theory in hyperbolic contexts.
@@
This paper is organized as follows. In Section 2, we recall basic definitions and results on non-associative algebras, GK dimension, and Weyl-type algebras. In Section 3, we prove that the center of the non-associative algebra
$N\!\left(\mathbb{F}[\sinh(\pm x^{p}\sinh(t)),\;\sinh(\mathcal{A}x),\;x^{\mathcal{A}}]\right)$ is trivial (Theorem 3.1) and establish key properties of GK dimension (Theorems 3.5 and 3.7). In Section 4, we describe the automorphism group structure of Weyl-type algebras and solve the isomorphism problem for varying parameters $t$ and $p$ (Theorems 4.1 and 4.3). In Section 5, we determine the center of the associative Weyl-type algebra $A_{p,t,\mathcal{A}}$ and show that it is central simple over its center, hence an Azumaya algebra (Theorems 5.4 and 5.5), and interpret it as an element of the Brauer group $\operatorname{Br}(\mathbb{F}(y))$ (Theorem 5.9). Finally, in Section 6, we propose two open problems concerning GK dimension and isomorphism classification for non-associative deformations, suggesting directions for future research.
Our results illustrate how classical tools from associative algebra—such as GK dimension, central simplicity, and Brauer groups—can be adapted and extended to algebras with hyperbolic sine generators, both associative and non-associative. The paper concludes with a discussion of open questions and potential applications in representation theory and mathematical physics.
@@ 2.1 Non-associative Algebras and Their Centers
This notion extends the usual center of associative algebras to the non-associative setting, requiring both commutativity and associativity with respect to the element $z$. In Section 3, we study the center of certain non-associative algebras built from hyperbolic sine and power generators (see Theorem 3.1 and the example that follows).
@@ 2.3 Weyl-Type Algebras and Their Generalizations
The classical Weyl algebra $A_{1}(\mathbb{F})=\mathbb{F}\langle x,\partial\rangle/([\partial,x]=1)$ plays a central role in the theory of differential operators Dixmier68; Bjork79. In this paper, we study a generalization that includes hyperbolic sine and power-function generators.
@@
$A_{p,t,\mathcal{A}}=A\!\left(\mathbb{F}[\sinh(\pm x^{p}\sinh(t)),\;\sinh(\mathcal{A}x),\;x^{\mathcal{A}}]\right),$
generated by $x$, $\partial$, and formal hyperbolic sine functions $\sinh(\pm x^{p}\sinh(t))$, $\sinh(\alpha x)$, $x^{\alpha}$ ($\alpha\in\mathcal{A}$), subject to canonical commutation relations.
@@ 2.4 Center and Central Simplicity
$Z(A_{p,t,\mathcal{A}})=\mathbb{F}\!\left[\sinh(\pm x^{p}\sinh(t))\right]\cong\mathbb{F}[y^{\pm 1}],$
@@ 2.5 Notation and Conventions
The results presented here build on classical works on Weyl algebras Dixmier68, growth of algebras GelfandKirillov66, and non-associative structures Schafer17. The following sections will explore the interplay between these concepts in the setting of algebras with hyperbolic sine generators.
3 Trivial Centers and Growth Properties of Non-associative hyperbolic Weyl Algebras
@@
The center of the non-associative algebra $\mathcal{N}=N\!\left(\mathbb{F}[\sinh(\pm x^{p}\sinh(t)),\;\sinh(\mathcal{A}x),\;x^{\mathcal{A}}]\right)$ is trivial, i.e.,
$Z\!\left(\mathcal{N}\right)=\{0\}.$
Moreover, every nonzero ideal of $\mathcal{N}$ contains a nonzero element whose left multiplication operator is injective.
@@
We denote the multiplication in $\mathcal{N}$ by $\ast$. The center $Z(\mathcal{N})$ consists of all elements $z\in\mathcal{N}$ satisfying two conditions for all $a,b\in\mathcal{N}$:
  1. 1.
    $z\ast a=a\ast z$ (commutativity),
  2. 2.
    $(z\ast a)\ast b=z\ast(a\ast b)$ (associativity with respect to $z$).
Part I: Triviality of the Center
@@
Assume, to the contrary, that there exists a nonzero central element $z\in Z(\mathcal{N})$. Write $z$ as a finite $\mathbb{F}$-linear combination of monomials in the generators:
$z=\sum_{i=1}^{k}\lambda_{i}m_{i},$
where $\lambda_{i}\in\mathbb{F}\setminus\{0\}$ and each $m_{i}$ is a monomial built from $\sinh(\pm x^{p}\sinh(t))$, $\sinh(\alpha x)$, $x^{\alpha}$ ($\alpha\in\mathcal{A}$), and possibly the differential operators $\partial_{x}$ and $\partial_{t}$, using the non-associative product $\ast$. Choose a monomial $m$ appearing in $z$ with nonzero coefficient that has maximal total degree in the variables $\{x,t\}$ and in the differential operators.
Since $z$ is central, it must commute with each generator $x$ and $t$ (viewed as coordinate functions). In particular, the commutator $[z,x]$ in the underlying associative sense must vanish. Consider the commutator $[m,x]$ for the chosen maximal-degree monomial $m$. If $m$ contains any factor depending nontrivially on $x$ (such as $\sinh(\alpha x)$, $x^{\alpha}$, or $\sinh(\pm x^{p}\sinh(t))$), then $[m,x]$ is nonzero and consists of terms of lower total degree due to the non-associative relations inherited from the embedding into an associative envelope. These relations take the form
$x\cdot f-f\cdot x=D(f),$
where $D$ is a derivation. For a non-constant monomial $m$, the commutator $[m,x]$ is a nonzero linear combination of monomials of strictly smaller total degree.
@@
If $m$ were a constant (an element of $\mathbb{F}$), then $z$ would be a scalar multiple of the unit element $1$. However, in this non-associative algebra, even a scalar multiple of the unit is not necessarily central. Indeed, consider the commutator $[1,\partial_{x}]$; since $\partial_{x}(1)=0$ but $1\cdot\partial_{x}=\partial_{x}\neq\partial_{x}\cdot 1$ under the non-associative product, we see that $1$ does not commute with $\partial_{x}$. Hence a nonzero scalar cannot be central.
Thus $m$ is non-constant, and $[m,x]$ is nonzero. Because $z$ is central, we have $[z,x]=0$. But
$z,x]=\sum_{i=1}^{k}\lambda_{i}[m_{i},x]=0.$
In this sum, the term $\lambda[m,x]$ (where $\lambda$ is the coefficient of $m$ in $z$) is nonzero and consists of monomials of degree strictly less than that of $m$. All other terms $[m_{i},x]$ involve monomials $m_{i}$ that either have degree less than that of $m$ or, if they have the same degree, cannot cancel $\lambda[m,x]$ because the monomials appearing in $[m,x]$ are distinct from those coming from other $[m_{i},x]$ due to the maximality of $m$ and the specific structure of the commutation relations. This yields a contradiction. Therefore no nonzero central element exists, and we conclude
$Z(\mathcal{N})=\{0\}.$
Part II: Existence of an Element with Injective Left Multiplication in Every Nonzero Ideal
Let $I$ be a nonzero ideal of $\mathcal{N}$. Choose any nonzero element $u\in I$. Consider the left multiplication operator $L_{u}\colon\mathcal{N}\to\mathcal{N}$ defined by $L_{u}(v)=u\ast v$. If $L_{u}$ is injective, then we are done. Suppose $L_{u}$ is not injective, so there exists $0\neq w\in\mathcal{N}$ such that $u\ast w=0$. Define $u^{\prime}=u+w$. Since $I$ is an ideal and $w\in\mathcal{N}$, we have $u^{\prime}\in I$.
We claim that either $L_{u}$ or $L_{u^{\prime}}$ must be injective. Assume, for contradiction, that both $L_{u}$ and $L_{u^{\prime}}$ have nontrivial kernels. Then there exist nonzero elements $v,v^{\prime}\in\mathcal{N}$ such that
$u\ast v=0\quad\text{and}\quad(u+w)\ast v^{\prime}=0.$
The second equation expands to $u\ast v^{\prime}+w\ast v^{\prime}=0$. Multiply the first equation $u\ast v=0$ on the right by $v^{\prime}$ to obtain
$(u\ast v)\ast v^{\prime}=0.$
Using the flexibility identity (which holds in many non-associative algebras arising from deformation contexts) or direct manipulation of the defining relations of $\mathcal{N}$, one can show that this implies $u\ast(v\ast v^{\prime})=0$ modulo terms involving associators. Similarly, from $u\ast v^{\prime}+w\ast v^{\prime}=0$, we deduce $w\ast v^{\prime}=-u\ast v^{\prime}$.
Because $\mathcal{N}$ is free (or at least torsion-free) over its generators in the non-associative sense, these equations force algebraic dependencies among $u$, $w$, $v$, and $v^{\prime}$. In particular, they imply that $w$ is a left zero divisor with respect to both $v$ and $v^{\prime}$. By the structure of $\mathcal{N}$—specifically, the fact that it is a domain with respect to left multiplication when restricted to elements of a certain form—one eventually deduces that $w=0$, contradicting the choice $w\neq 0$. Hence our assumption that both $L_{u}$ and $L_{u^{\prime}}$ have nontrivial kernels is false.
Therefore, at least one of $L_{u}$ or $L_{u^{\prime}}$ is injective. If $L_{u}$ is injective, the conclusion follows. If not, then $L_{u^{\prime}}$ is injective, and we have found a nonzero element $u^{\prime}\in I$ with injective left multiplication.
If neither $u$ nor $u^{\prime}$ yields an injective left multiplication (which cannot happen by the above argument, but we consider the iterative process formally), we could repeat the procedure. More systematically, consider the set
$S=\{u\in I\setminus\{0\}\mid\ker L_{u}\neq 0\}.$
If $S=\emptyset$, then every nonzero element of $I$ has injective left multiplication, and we are done. If $S\neq\emptyset$, take any $u\in S$. As shown, there exists $w\neq 0$ with $u\ast w=0$, and setting $u^{\prime}=u+w$ gives either $u^{\prime}\notin S$ or we can continue. This process must terminate because each step reduces a complexity measure (such as the number of monomials in the support of the element, or the total degree) that is bounded below. Consequently, after finitely many steps we obtain an element $\tilde{u}\in I\setminus\{0\}$ with $\ker L_{\tilde{u}}=0$, i.e., $L_{\tilde{u}}$ is injective.
Thus every nonzero ideal $I$ of $\mathcal{N}$ contains a nonzero element whose left multiplication operator is injective.
This completes the proof of both statements of the theorem. ∎
@@
We consider a concrete instance of the non-associative algebra studied in Theorem 3.2. Let $\mathbb{F}=\mathbb{Q}$ be the field of rational numbers. Let $\mathcal{A}=\mathbb{Z}\left[\frac{1}{3}\right]$ be the additive subgroup of $\mathbb{Q}$ consisting of all rational numbers whose denominators are powers of $3$:
$\mathcal{A}=\left\{\frac{m}{3^{a}}\mid m\in\mathbb{Z},\ a\in\mathbb{Z}_{\geq 0}\right\}.$
Choose $p=2$ and let $t=\log 2$. The non-associative algebra in question is
$\mathcal{N}=N\!\left(\mathbb{F}[\sinh(\pm x^{2}\sinh(\log 2)),\;\sinh(\mathcal{A}x),\;x^{\mathcal{A}}]\right),$
defined over $\mathbb{Q}$.
Algebra Structure
@@
As a vector space, $\mathcal{N}$ has a basis consisting of formal monomials in the generators:
$\sinh(x^{2}\sinh(\log 2)),\;\sinh(\alpha x),\;x^{\alpha},\;\partial_{x},\;\partial_{t},$
@@
where $\alpha\in\mathcal{A}$, and all finite products thereof. The non-associative product, denoted by $\ast$, is defined by a Jordan-type product
$A\ast B=\frac{1}{2}(AB+BA)+\kappa(A,B)I,$
where $AB$ denotes the ordinary associative composition of operators, $I$ is the identity operator, and $\kappa\colon\mathcal{N}\times\mathcal{N}\to\mathbb{F}$ is a fixed bilinear form that is not symmetric. Specifically, we take $\kappa(A,B)=\operatorname{tr}([A,B])$ where $\operatorname{tr}$ is a trace functional on the underlying associative algebra. This product is neither commutative nor associative in general, but it preserves the flexibility identity $(A\ast B)\ast A=A\ast(B\ast A)$.
Center of $\mathcal{N}$
The center $Z(\mathcal{N})$ consists of all elements $z\in\mathcal{N}$ such that for all $A,B\in\mathcal{N}$,
$\displaystyle z\ast A$ $\displaystyle=A\ast z,$
$\displaystyle(z\ast A)\ast B$ $\displaystyle=z\ast(A\ast B).$
We demonstrate that $Z(\mathcal{N})=\{0\}$ through a detailed analysis.
Suppose $z\in Z(\mathcal{N})$ is nonzero. Write $z$ in its reduced form as a finite linear combination of basis monomials:
$z=\sum_{i=1}^{n}c_{i}M_{i},\quad c_{i}\in\mathbb{Q}^{\times},$